OpenStudy (anonymous):

Which is the graph of the line with equation: y + 4 = 2(x – 1)

OpenStudy (anonymous):

I think it's A

OpenStudy (cicilybailey14):

C I think

OpenStudy (phi):

start with y + 4 = 2(x – 1) test a few x numbers. example. if x is 1, y+4= 2(1-1) y+4= 0 y= -4